OpenStudy (anonymous):

If I start out making $15 dollars a day, and my pay is increased by $0.75 a day, how much money will I have at the end of the month? (31 days)?

OpenStudy (anonymous):

It's 0.75 ADDED to the INCREASED amount! Which means I start out with 15, then on to 15.75, then on to 16.50, and so on!

OpenStudy (anonymous):

15,15.75,16.5,... An=a1+r.(n-1) Sn=(a1+an).n/2 The final answer is: 813,75

OpenStudy (anonymous):

Whoa... are you sure??

OpenStudy (anonymous):

Yes. Do you not have the answer?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

I do It by arithmetical progression. I would say that this is right, but i'm not a teacher, hehehe.

OpenStudy (anonymous):

OOOH wait, is that supposed to be a decimal point?

OpenStudy (anonymous):

Oh sorry. I misread it. You're right.

OpenStudy (anonymous):

Wait shouldn't it be 922.25?

OpenStudy (anonymous):

The final answer is: 813.75 Yes, It is a decimal point, i'm brazilian and we use "," in place of ".".

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

I added 15 increasing 1, 31 times lol http://www.wolframalpha.com/input/?i=15%2B16%2B17%2B18%2B19%2B20%2B21%2B22%2B23%2B24%2B25%2B26%2B27%2B28%2B29%2B30%2B31%2B32%2B33%2B34%2B35%2B36%2B37%2B38%2B39%2B40%2B41%2B42%2B43%2B44%2B45 Then I multiplied .75 by 31, which is 23.25, then I subtracted that by the total amount of 930, which is 906 (i messed up)

OpenStudy (anonymous):

Or am I completely wrong? I admit I have no idea what I'm doing

OpenStudy (anonymous):

I multiplied .75 by 31 to deduct it from the dollar

OpenStudy (anonymous):

No, It's like the compound interest. The first is 15, the second ir 15,75, the third is 16,50.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

It's a arithmetical progression.